# ReSem3D: Refinable 3D Spatial Constraints via Fine-Grained Semantic Grounding for Generalizable Robotic Manipulation

### v3

@@ III-B1 Part-Level Constraint Extraction
Step 2: Part-Level Semantic Consistency Clustering. Density-based clustering (DBSCAN) [ref43] is performed on the filtered mask set to further reduce visual ambiguity and redundancy, with masks of identical semantic labels merged within each cluster. The clustered mask set is denoted as
@@
where $M_{\text{cluster}}=\{{\widetilde{m}}_{1},{\widetilde{m}}_{2},\ldots,{\widetilde{m}}_{e}\}$, and $e$ is the number of clustered masks. To further extract effective part-level constraints, we define the centroid coordinate $\bm{c}_{j}$ for each clustered mask ${\widetilde{m}}_{j}\in M_{\text{cluster}}$ as
@@
where $(x,y)$ denotes the 2D coordinates of the pixels in the mask $\widetilde{m}_{j}$, and $|\widetilde{m}_{j}|$ is the number of pixels in $\widetilde{m}_{j}$.
Step 3: Part-Level Constraint Extraction. Each centroid in the coordinate set $C=\{\bm{c}_{1},\bm{c}_{2},\ldots,\bm{c}_{q}\}$ is sequentially assigned a numeric label from the label set $\mathcal{Z}=\{0,1,\ldots,q-1\}$ to enhance visual grounding of MLLM. These labels are overlaid onto the original RGB image $I_{\text{rgb}}\in\mathbb{R}^{H\times W\times 3}$ to form an explicit visual prompt. Together with the subtask instruction $L_{i}$, the resulting multimodel input is fed into the MLLM, which leverages semantic grounding to extract the labels relevant to the instruction. Finally, the extracted labels are projected into 3D space via the depth information $I_{\text{depth}}\in\mathbb{R}^{H\times W}$, completing the localization of initial constraint $\mathcal{C}_{i}^{\text{init}}$.
III-B2 Region-Level Constraint Refinement
As shown in Fig. 3, to enhance the granularity and adaptability of spatial constraints, we propose a semantic-adaptive region-level refinement method that automatically extracts semantic information and selects between two refinement strategies: geometric constraint refinement and positional constraint refinement. The former focuses on fine-grained structural priors within the mask region, while the latter emphasizes the spatial layout of the semantically mapped area.
Geometric Constraint Refinement: Given the mask ${\widetilde{m}}_{j}$ associated with the extracted centroid coordinate $\bm{c}_{j}\in C$ by the MLLM, the geometric constraint refinement is applied to enhance fine-grained visual grounding and proceeds through the following five steps.
Figure 3: Region-Level Constraint Refinement. Within the MLLM-driven TAMP framework, this module encompasses two strategies: geometric and positional constraint refinement. The geometric refinement adjusts the tweezers’ grasp point from the center to the two tips, thereby introducing detailed geometric priors. The positional refinement localizes the trash bin’s placement to the midpoint between symmetrical centers, thereby enhancing the spatial precision.
Step 1: Mask Normalization. Proportional discrepancies among different masks ${\widetilde{m}}_{j}$ are mitigated by cropping the effective region of each mask and uniformly scaling them, thereby embedding all masks consistently onto the fixed-resolution standard canvas. Specifically, we first extract coordinates of the foreground pixels in ${\widetilde{m}}_{j}$ as
@@ III-B2 Region-Level Constraint Refinement
Step 2: Grid Construction and Label Embedding. In order to improve visual grounding of geometric structures, grid construction and label embedding are performed on the target mask image. Specifically, the $j$-th mask ${\widetilde{m}}_{j}^{t}$ is divided into $G_{h}\times G_{w}$ grids, where the width $\omega_{c}$ and height $h_{c}$ of each grid are determined by the target resolution $H_{t}\times W_{t}$ as follows:
@@
where $\mathcal{P}_{k,s}^{j}=\sum_{(x,y)\in\mathcal{G}_{k,s}}\left[\tilde{m}_{j}^{t}(x,y)=1\right]$ denotes the set of foreground pixels within $\mathcal{G}_{k,s}^{j}$, and $|\mathcal{P}_{k,s}^{j}|$ is the number of pixels. When the foreground density $\rho_{k,s}^{j}$ surpasses the assignment threshold $\tau$, the semantic label $\ell_{k,s}^{j}\in\left\{0,\ 1,\ldots,N_{\text{valid}}-1\right\}$ is assigned to the grid. We define the mapping $\mathcal{U}:\ \mathcal{G}_{k,s}^{j}\rightarrow\ell_{k,s}^{j}$ to associate each grid with the semantic label, where $N_{valid}$ indicates the total number of grids satisfying the threshold condition. The centroid within the grid are then calculated as
@@
Step 3: Centroid mapping and label association. The centroid coordinates $\bm{c}_{j}\in C$ extracted from the part-level constraints are mapped onto the target mask and associated with the grid semantic labels $\ell_{k,s}^{j}$ , which are overlaid on the target mask as textual prompts to enhance the fine-grained geometric understanding of MLLM. Subsequently, we map the centroid $\bm{c}_{j}=\begin{bmatrix}x_{j}\\ y_{j}\end{bmatrix}$ into the the target mask as follows:
@@
where $\bar{\bm{c}}_{j}=\begin{bmatrix}\bar{x}_{j}\\ \bar{y}_{j}\end{bmatrix}$ denotes the coordinate in the target mask. To establish the correspondence with grid labels, $\bar{\bm{c}}_{j}$ is assigned to its corresponding grid $\mathcal{G}_{k,s}^{j}$ by Equation (15), and then projected to the grid centroid $\bm{c}_{j}^{k,s}$ following Equation (17), thereby associating it to the semantic label $\ell_{k,s}^{j}$.
Step 4: Fine-grained semantic reasoning by MLLM. Achieving region-level constraint refinement and deriving new labels ${\hat{\ell}}_{k,s}^{j}$ based on the target geometric involve feeding the original RGB image $I_{\text{rgb}}$ , the original mask ${\widetilde{m}}_{j}$ , and the target mask ${\widetilde{m}}_{j}$ as visual prompts into the MLLM. Subsequently, the numerical label $z_{j}$ and the semantic label $\ell_{k,s}^{j}$ are presented as textual prompts to the MLLM for reasoning.
Step 5: Region-Level Constraint Extraction. The refined constaint is constructed by inversely mapping the labels ${\hat{\ell}}_{k,s}^{j}$ to their corresponding grids $\mathcal{G}_{k,s}^{j}$ via the instantiated mapping $\mathcal{U}:\mathcal{G}_{k,s}^{j}\rightarrow\ell_{k,s}^{j}$. The associated grid centroids $\bm{c}_{j}^{k,s}$ are then projected to the original mask coordinate as $\bm{c}_{j}$ using Equations (17) and (18). Finally, the 2D coordinates are transformed into 3D space under the depth information $I_{\text{depth}}\in\mathbb{R}^{H\times W}$, thereby forming the initial constraint $\mathcal{C}_{i}^{\text{init}}$.
Positional Constraint Refinement: To address positional discrepancies in semantic grounding of objects characterized by distinct boundary features, including open-top structures, we propose a three-step refinement method that exploits edge points and spatial geometry. By integrating density peak estimation with symmetry analysis, this approach effectively mitigates MLLM limitations in position modeling, enhancing the practical feasibility of semantic grounding.
Step 1: Extraction of 3D Coordinates Along Mask Edges. The edge point set $\mathcal{M}_{j}$ associated with mask ${\widetilde{m}}_{j}^{t}$ is defined for extracting edge points of the object as follows:
$\mathcal{M}_{j}=\{\left(x,y\right)\ |\ {\widetilde{m}}_{j}^{t}(x,y)=1\}$ (19)
@@
$\exists(x\prime,y\prime)\in\mathcal{N}_{4}\left(x,y\right)\quad\text{s.t.}{\ \widetilde{m}}_{j}^{t}(x\prime,y\prime)=0$
@@
Step 2: Density Peak Estimation and Maximum Height Filtering. Kernel Density Estimation (KDE) is employed to model the $z$-axis distribution of edge points, thereby facilitating the elimination of outliers and extraction of prominent edge features. Given the set $\{z_{i}\}_{i=1}^{N}$, which represents the heights of $N$ edge points obtained from the mask denoted by ${\widetilde{m}}_{j}^{t}$, the probability density function is approximated by
@@
where $K\left(\cdot\right)$ denotes the Gaussian kernel function and $h$ is the bandwidth. The density peak is defined as $z^{\ast}={\arg\max}_{z}\hat{f}(z)$. All points within the height range $\left|z_{i}-z^{\ast}\right|\leq\delta$ are retained, resulting in filtered 3D and 2D point sets $\mathcal{P}_{j}^{\text{KDE}}\subset\mathcal{P}_{j}$ and $\mathcal{M}_{j}^{\text{KDE}}\subset\mathcal{M}_{j}$, respectively. Subsequently, the maximum height along the $z$-axis in $\mathcal{P}_{j}^{\text{KDE}}$ is identified as $z_{max}$. Finally, points with heights satisfying $z_{i}{\geq z}_{max}-\eta$ are further selected, yielding the final refined point sets $\mathcal{P}_{j}^{\text{peak}}\subset\mathcal{P}_{j}^{\text{KDE}}$ and $\mathcal{M}_{j}^{\text{peak}}\subset\mathcal{M}_{j}^{\text{KDE}}$.
Step 3: Extraction of centrally symmetric point pairs. Further integration of object position and structural information is achieved by identifying centrally symmetric point pairs in 2D space, which are used to derive the refined 3D spatial constraints, denoted as $\mathcal{C}_{i}^{\text{init}}$. Given the centroid $\bm{c}_{j}$ , we enumerate all point pairs $\bm{u}_{a},\bm{u}_{b}\in\mathcal{M}_{j}^{\text{peak}}$ and identify the pair for which the midpoint is closest to $\bm{c}_{j}$, which is defined as
@@
where $\left(\bm{u}_{a}^{\ast},\bm{u}_{b}^{\ast}\right)$ denotes the filtered point pair, with 3D coordinates $\bm{p}_{a},\bm{p}_{b}\in\mathcal{P}_{j}^{\text{peak}}$. Finally, the refined 3D spatial constraint $\mathcal{C}_{i}^{\text{init}}$ is expressed as
@@ III-C Real-Time Close-Loop Control Strategy
To solve the constrained optimization problem in Equation (2) with real-time performance, the Model Predictive Path Integral (MPPI) control algorithm is employed on the Isaac Gym simulation platform. MPPI is a stochastic optimization method designed for robot control in dynamic and uncertain environments. Essentially, the algorithm samples multiple candidate control sequences from the control input space, evaluates the cost functions based on simulated state trajectories, and updates the control policy through importance-weighted averaging. The GPU-accelerated parallel simulation enabled by Isaac Gym significantly enhances sampling efficiency, thereby supporting real-time execution in dynamic environments. For a detailed explanation of MPPI, please refer to [ref44, ref45]. In this work, we derive the specific formulation for solving the constrained optimization problem presented in Equation (2).
@@
where $\bm{x}_{k}(t+1)=S(\bm{x}_{k}(t),\bm{v}_{k}(t))$. To further solve the problem, we introduce the forward kinematics mapping $\varphi:\ \mathbb{R}^{2n}\rightarrow SE(3)$, which maps the joint space state to the end-effector pose, expressed as $\bm{T}_{e}^{k}(t)=\varphi(\bm{x}_{k}(t))$. At this point, $\bm{T}_{e}^{k}(t)\in SE(3)$ denotes the end-effector pose of the $k$-th trajectory at time $t$. For each trajectory $\mathcal{X}_{k}$, the total cost is the cumulative sum of the objective function $\mathcal{J}\left(\cdot\right)$ over each time step, defined as
@@
where $P(\cdot)$ and $R(\cdot)$ denote the position and quaternion components extracted from the spatial pose. The errors are calculated as $d_{p}(\bm{p}_{1},\bm{p}_{2})={\left\|\bm{p}_{1}-\bm{p}_{2}\right\|}_{2}$ and $d_{r}(\bm{q}_{1},\bm{q}_{2})=2\,\text{arccos}(\text{max}(-1,\text{min}(1,\langle\bm{q}_{1},\bm{q}_{2}\rangle)))$. Subsequently, the objective function is defined as
@@
where $\lambda_{p}$ and $\lambda_{r}$ are the weights for position and orientation costs, $\lambda_{c}$ denote the weight for the collision cost $\mathcal{D}_{c}$, which is formulated based on the ESDF representation, as described in [ref9]. In this work, the objective function is subject to preconditions and postconditions as specified in Equation (2), the precondition typically ensures that the end-effector is within the designated region prior to the execution stage, while the postcondition assesses whether the end-effector has successfully reached the desired pose, defined as

### v4

@@ III-B1 Semantic Structuring Representation
Step 2: Visual Consistency Clustering. Density-based clustering (DBSCAN) is performed on the filtered mask set to further reduce visual ambiguity and redundancy, with masks of identical semantic labels merged within each cluster. The clustered mask set is denoted as
@@
where $M_{\text{cluster}}=\{{\widetilde{m}}_{1},{\widetilde{m}}_{2},\ldots,{\widetilde{m}}_{e}\}$, and $e$ is the cardinality.
Coarse Affordance Anchor. Since affordance targets requires reasoning within affordance regions while remaining consistent with implicit semantic intent, we introduce a coarse affordance anchor to instantiate an optimizable starting point for downstream refinement. We define the centroid coordinate $\bm{c}_{j}$ for each clustered mask ${\widetilde{m}}_{j}\in M_{\text{cluster}}$ as
@@
where $(x,y)$ denotes the 2D coordinates of the pixels in the mask $\widetilde{m}_{j}$, and $|\widetilde{m}_{j}|$ is the number of pixels in $\widetilde{m}_{j}$. Each centroid in the coordinate set $C=\{\bm{c}_{1},\bm{c}_{2},\ldots,\bm{c}_{q}\}$ is sequentially assigned a numeric label from the label set $\mathcal{Z}=\{0,1,\ldots,q-1\}$ to enhance semantic grounding of MLLM. These labels are overlaid onto the original RGB image $I_{\text{rgb}}\in\mathbb{R}^{H\times W\times 3}$ to form an explicit visual prompt. Together with the subtask instruction $L_{i}$, the resulting multimodel input is fed into the MLLM, which leverages semantic grounding to extract the labels relevant to the instruction, thereby obtaining the affordance anchor $C_{\text{anchor}}\subseteq C$. The $M_{\text{region}}\subseteq M_{\text{cluster}}$ associated with $C_{\text{anchor}}$ is constructed as the affordance region. Finally, the extracted labels are projected into 3D space via the depth information $I_{\text{depth}}\in\mathbb{R}^{H\times W}$, completing the instantiation of $C_{\text{anchor}}$. Based on the above formulation, the semantic affordance description is represented as $\mathcal{A}=\{S_{\text{intent}},M_{\text{region}},C_{\text{anchor}}\}$.
III-B2 Affordance Refinement Strategy
As shown in Fig. 3, to effectively leverage the semantic affordance description $\mathcal{A}$ and further refine the coarse affordance anchor $C_{\text{anchor}}$, we propose an Affordance Refinement Strategy driven by the MLLM. This strategy automatically parses $\mathcal{A}$ and selects between two refinement flows: a geometric refinement flow, which focuses on object-level geometric structure, and a positional refinement flow, which emphasizes object opening localization. In addition, the instantiated $C_{\text{anchor}}$ will be directly used as the affordance targets $\mathcal{C}_{i}^{\text{init}}$ when the implicit intent cannot be resolved by the MLLM.
Geometric Refinement Flow: Given the semantic affordance description $\mathcal{A}=\{S_{\text{intent}},M_{\text{region}},C_{\text{anchor}}\}$, the geometric refinement flow is applied to enhance fine-grained semantic grounding and proceeds through the following five steps.
Figure 3: Affordance Refinement Strategy. Within the MLLM-driven TAMP framework, this module encompasses two strategies: geometric and positional refinement. The geometric refinement adjusts the tweezers’ grasp point from the center to the two tips, thereby introducing detailed geometric priors. The positional refinement localizes the trash bin’s placement to the midpoint between symmetrical centers, thereby enhancing the spatial precision.
Step 1: Mask Normalization. Proportional discrepancies among different masks ${\widetilde{m}}_{j}\in M_{\text{region}}$ are mitigated by cropping the effective region of each mask and uniformly scaling them, thereby embedding all masks consistently onto the fixed-resolution standard canvas. Specifically, we first extract coordinates of the foreground pixels in ${\widetilde{m}}_{j}$ as
@@ III-B2 Affordance Refinement Strategy
Step 2: Grid Construction and Label Embedding. In order to improve semantic grounding of geometric structures, grid construction and label embedding are performed on the target mask. Specifically, the $j$-th mask ${\widetilde{m}}_{j}^{t}$ is divided into $G_{h}\times G_{w}$ grids, where the width $\omega_{c}$ and height $h_{c}$ of each grid are determined by the target resolution $H_{t}\times W_{t}$ as follows:
@@
where $\mathcal{P}_{k,s}^{j}=\sum_{(x,y)\in\mathcal{G}_{k,s}}\left[\tilde{m}_{j}^{t}(x,y)=1\right]$ denotes the set of foreground pixels within $\mathcal{G}_{k,s}^{j}$, and $|\mathcal{P}_{k,s}^{j}|$ is the number of pixels. When the foreground density $\rho_{k,s}^{j}$ surpasses the assignment threshold $\tau$, the semantic label $\ell_{k,s}^{j}\in\left\{0,\ 1,\ldots,N_{\text{valid}}-1\right\}$ is assigned to the grid. We define the mapping $\mathcal{U}:\ \mathcal{G}_{k,s}^{j}\rightarrow\ell_{k,s}^{j}$ to associate each grid with the semantic label, where $N_{\text{valid}}$ indicates the total number of grids satisfying the threshold condition. The centroid within the grid are then calculated as
@@
Step 3: Centroid mapping and label association. The course anchor $\bm{c}_{j}\in C_{\text{anchor}}$ are mapped onto the target mask and associated with the grid semantic labels $\ell_{k,s}^{j}$ , which are overlaid on the target mask as textual prompts to enhance the fine-grained geometric understanding of MLLM. Subsequently, we map the anchor $\bm{c}_{j}=[x_{j}\;y_{j}]^{\top}$ into the the target mask as follows:
@@
where $\bar{\bm{c}}_{j}=[\bar{x}_{j}\;\bar{y}_{j}]^{\top}$ denotes the anchor in the target mask. To establish the correspondence with grid labels, $\bar{\bm{c}}_{j}$ is assigned to its corresponding grid $\mathcal{G}_{k,s}^{j}$ by Equation (15), and then projected to the grid centroid $\bm{c}_{j}^{k,s}$ following Equation (17), thereby associating it to the semantic label $\ell_{k,s}^{j}$.
Step 4: Fine-grained Geometric Reasoning by MLLM. To generate new labels ${\hat{\ell}}_{k,s}^{j}$ based on the geometric structure, we feed the original RGB image $I_{\text{rgb}}$, the original mask ${\widetilde{m}}_{j}$, and the target mask ${\widetilde{m}}_{j}^{t}$ as visual prompts into the MLLM. Subsequently, the numerical label $z_{j}\in\mathcal{Z}$, the semantic label $\ell_{k,s}^{j}$, and the implicit intent $s_{\text{g}}\in S_{\text{intent}}$ are provided as textual prompts to the MLLM for reasoning.
Step 5: Refined Anchors as Affordance Targets. Through the instantiated mapping $\mathcal{U}:\mathcal{G}_{k,s}^{j}\rightarrow\ell_{k,s}^{j}$, the newly generated labels ${\hat{\ell}}_{k,s}^{j}$ are inversely mapped to their corresponding grids $\mathcal{G}_{k,s}^{j}$, thereby constructing the refined anchor. Subsequently, the associated grid centroids $\bm{c}_{j}^{k,s}$ are projected to the original mask coordinate as $\bm{c}_{j}$. Finally, the 2D coordinates are transformed into 3D space under the depth information $I_{\text{depth}}\in\mathbb{R}^{H\times W}$, forming the affordance targets $\mathcal{C}_{i}^{\text{init}}$.
Positional Refinement Flow: To mitigate positional discrepancies in semantic grounding of objects characterized by distinct boundary features, including open-top structures, we propose a three-step refinement method that exploits edge points and affordance regions to optimize $\mathcal{A}=\{S_{\text{intent}},M_{\text{region}},C_{\text{anchor}}\}$. By integrating density peak estimation with symmetry analysis, this approach effectively mitigates MLLM limitations in position modeling, enhancing the practical feasibility of semantic grounding.
Step 1: Extraction of 3D Coordinates Along Mask Edges. The edge point set $\mathcal{M}_{j}$ associated with mask ${\widetilde{m}}_{j}\in M_{\text{region}}$ is defined for extracting edge points of the object as follows:
$\mathcal{M}_{j}=\{\left(x,y\right)\ |\ {\widetilde{m}}_{j}(x,y)=1\}$ (19)
@@
$\exists(x\prime,y\prime)\in\mathcal{N}_{4}\left(x,y\right)\quad\text{s.t.}{\ \widetilde{m}}_{j}(x\prime,y\prime)=0$
@@
Step 2: Density Peak Estimation and Maximum Height Filtering. Kernel Density Estimation (KDE) is employed to model the $z$-axis distribution of edge points, thereby facilitating the elimination of outliers and extraction of prominent edge features. Given the set $\{z_{i}\}_{i=1}^{N}$, which represents the heights of $N$ edge points obtained from the mask denoted by ${\widetilde{m}}_{j}$, the probability density function is approximated by
@@
where $K\left(\cdot\right)$ denotes the Gaussian kernel function and $h$ is the bandwidth. The density peak is defined as $z^{\ast}={\arg\max}_{z}\hat{f}(z)$. All points within the height range $\left|z_{i}-z^{\ast}\right|\leq\delta$ are retained, resulting in filtered 3D and 2D point sets $\mathcal{P}_{j}^{\text{KDE}}\subset\mathcal{P}_{j}$ and $\mathcal{M}_{j}^{\text{KDE}}\subset\mathcal{M}_{j}$. Subsequently, the maximum height along the $z$-axis in $\mathcal{P}_{j}^{\text{KDE}}$ is identified as $z_{max}$. Finally, points with heights satisfying $z_{i}{\geq z}_{max}-\eta$ are further selected, yielding the final point sets $\mathcal{P}_{j}^{\text{peak}}\subset\mathcal{P}_{j}^{\text{KDE}}$ and $\mathcal{M}_{j}^{\text{peak}}\subset\mathcal{M}_{j}^{\text{KDE}}$.
Step 3: Centrally Symmetric Point Pairs as Affordance Targets. Further integration of object position and structural information is achieved by identifying centrally symmetric point pairs, which are used to derive the refined affordance targets, denoted as $\mathcal{C}_{i}^{\text{init}}$. Given the course anchor $\bm{c}_{j}\in C_{\text{anchor}}$, we enumerate all point pairs $\bm{u}_{a},\bm{u}_{b}\in\mathcal{M}_{j}^{\text{peak}}$ and identify the pair whose midpoint is closest to $\bm{c}_{j}$, defined as
@@
where $\left(\bm{u}_{a}^{\ast},\bm{u}_{b}^{\ast}\right)$ denotes the filtered point pair, with 3D coordinates $\bm{p}_{a},\bm{p}_{b}\in\mathcal{P}_{j}^{\text{peak}}$. Finally, the refined affordance target $\mathcal{C}_{i}^{\text{init}}$ is expressed as
@@ III-C Real-Time Close-Loop Control Strategy
To solve the constrained optimization problem in Equation (2) with real-time performance, the Model Predictive Path Integral (MPPI) control algorithm is employed on the Isaac Gym simulation platform. MPPI is a stochastic optimization method designed for robot control in dynamic and uncertain environments. Essentially, the algorithm samples multiple candidate control sequences from the control input space, evaluates the cost functions based on simulated state trajectories, and updates the control policy through importance-weighted averaging. The GPU-accelerated parallel simulation enabled by Isaac Gym significantly enhances sampling efficiency, thereby supporting real-time execution in dynamic environments. For a detailed explanation of MPPI, please refer to [williams2018mppi]. In this work, we derive the specific formulation for solving the constrained optimization problem presented in Equation (2).
@@
where $\bm{x}_{k}(t+1)=S(\bm{x}_{k}(t),\bm{v}_{k}(t))$. We further introduce the forward kinematics mapping $\varphi:\ \mathbb{R}^{2n}\rightarrow SE(3)$, which maps the joint space state to the end-effector pose, expressed as $\bm{T}_{e}^{k}(t)=\varphi(\bm{x}_{k}(t))$. At this point, $\bm{T}_{e}^{k}(t)\in SE(3)$ denotes the end-effector pose of the $k$-th trajectory at time $t$. For each trajectory $\mathcal{X}_{k}$, the total cost is the cumulative sum of the objective function $\mathcal{J}\left(\cdot\right)$ over each time step, defined as
@@
where $P(\cdot)$ and $R(\cdot)$ denote the position and quaternion components. The errors are calculated as $d_{p}(\bm{p}_{1},\bm{p}_{2})={\left\|\bm{p}_{1}-\bm{p}_{2}\right\|}_{2}$ and $d_{r}(\bm{q}_{1},\bm{q}_{2})=2\,\text{arccos}(\text{max}(-1,\text{min}(1,\langle\bm{q}_{1},\bm{q}_{2}\rangle)))$. Subsequently, the objective function is defined as
@@
where $\lambda_{p}$ and $\lambda_{r}$ are the weights for position and orientation costs, $\lambda_{c}$ denote the weight for the collision cost $\mathcal{D}_{c}$, which is formulated based on the ESDF representation, as described in [huang2024rekep]. In this work, the objective function is subject to preconditions and postconditions as specified in Equation (2), the precondition typically ensures that the end-effector is within the designated region prior to the execution stage, while the postcondition assesses whether the end-effector has successfully reached the desired pose, defined as